OpenStudy (anonymous):

If f(x) = 3x^2 - x + 2 find [f(a)]^2 please show work

OpenStudy (anonymous):

\[f(a) = 3a^2-x+2\]\[[ f(a) ]^2 = [ 3a^2-x+2 ] ^2 = [3a^2-x+2][3a^2-x+2]\] can you go from here?

OpenStudy (anonymous):

typo.. it's supose to be "a" for all x'es

OpenStudy (anonymous):

sorry no I can't

OpenStudy (anonymous):

\[[3a^2−a+2][3a^2−a+2] = (3a^2)(3a^2)+(3a^2)(-a)+(3a^2)(2)+(-a)(3a^2)+(-a)(-a)+(-a)(2)+(2)(3a^2)+(2)(-a)+(2)(2)\]

OpenStudy (anonymous):

it didn't fit here lol

OpenStudy (anonymous):

ok i think i understand now

OpenStudy (anonymous):

but basically its: \[[3a^2−a+2][3a^2−a+2] = \] \[(3a^2)(3a^2)+(3a^2)(-a)+(3a^2)(2)\]\[+(-a)(3a^2)+(-a)(-a)+(-a)(2) \]\[+(2)(3a^2)+(2)(-a)+(2)(2)=\] \[9a^4 - 3a^3 -3a^3 + 6a^2 -a^2+6a^2-2a-2a+4\] is that better?